# Sequential Attention-based Sampling for Histopathological Analysis

### v2

@@ A.5 Hardware
For feature extraction and model inference, we utilized an NVIDIA GeForce GTX 1080 Ti GPU with 12 GB of memory. Feature extraction was carried out with a batch size of 512. Training for HAFED and TSU was also performed on the same GPU, while to accelerate the training of the RL agent described in Section 3.4, we employed an NVIDIA Tesla V-100 GPU with 32 GB of memory.
A.6 Limitations
@@
Fixed Budget Allocation - In our current setup, a fixed sampling budget per slide (e.g., 10% or 20%) is used. However, as shown in Figure 4, we observe that for certain slides, the classifier becomes confident early in the sampling process, rendering additional samples redundant. This suggests that a dynamic budget allocation strategy—guided by the classifier’s confidence—could further reduce inference time without compromising performance.
[…]
@@
Selection of Normal Patches - WSI usually contains more normal patches compared to tumor patches. This imbalance becomes particularly problematic when the tumor occupies only a small fraction of the slide. In such cases, the RL agent is more likely to encounter and select normal patches—especially in the early steps—leading to inefficient exploration and increased risk of misclassification. The sparse distribution of tumor regions makes it difficult for the agent to consistently focus on diagnostically relevant areas within a limited sampling budget.A more efficient strategy would involve the early discarding of low-confidence or redundant regions, allowing the agent to concentrate its computational efforts on a smaller, more informative subset of patches. This could improve both performance and inference efficiency.

### v3

@@ A.5 Hardware
For feature extraction and model inference, we utilized an NVIDIA GeForce GTX 1080 Ti GPU with 12 GB of memory. Feature extraction was carried out with a batch size of 512. Training for HAFED and TSU was also performed on the same GPU. To accelerate the training of the RL agent described in Section˜3.4, we employed an NVIDIA Tesla V-100 GPU with 32 GB of memory.
A.6 Limitations and Extensions
Training time cost - A key limitation of our method is that, despite being RL-based, it performs feature extraction on all high-resolution patches during training. To reduce training time, future improvements would involve developing an end-to-end pipeline that trains the RL agent concurrently with the HAFED and the TSU modules. In such a model, the policy network decides the specific low-resolution patch to sample in real-time, followed by HAFED feature distillation for the zoomed-in patch, and TSU state update for the entire slide, with all steps being trained end-to-end with reward at each time step. This would be followed by the additional steps of training the policy and value networks, and so on, until the end of each episode, for several episodes. This mimics how RL agents learn, conventionally, in novel environments with sporadic rewards. With this method, there would be no need to sample every patch at high resolution during training. Yet, key challenges – with stable, training of multiple modules concurrently – may need to be addressed.
@@ A.6 Limitations and Extensions
Fixed sampling budget - In our current setup, we employ a fixed sampling budget per slide (e.g., 10% or 20%). However, as shown in Figure˜4, we observe that for certain slides, the classifier can become confident in its prediction fairly early in the sampling process, rendering further sampling unnecessary. This suggests that a dynamic budget allocation strategy – guided by the classifier’s confidence – could further reduce inference time without compromising performance.
Architecture overhead and parameter scaling - We compare the total number of model parameters—excluding the ViT-based feature extractor (see Section˜A.2.1)—across different methods. DTFD, ACMIL, and HAFED, all of which operate on high-resolution inputs, are relatively lightweight with approximately $82.56\text{K}$, $84.37\text{K}$, and $300\text{K}$ parameters, respectively. In contrast, the SASHA pipeline consists of three distinct modules: HAFED, the Target State Updater (TSU), and the Reinforcement Learning (RL) agent, with parameter counts of $300\text{K}$, $1.1\text{M}$, and $296\text{K}$, respectively—summing to approximately $1.7\text{M}$ parameters. This represents a $3.5\times$ increase in parameters compared to HAFED and a $21\times$ increase relative to DTFD and ACMIL.
@@
Filtering normal patches - WSIs typically contain far more normal patches, as compared to tumor patches. This imbalance becomes particularly problematic when the tumor occupies only a small fraction of the slide. In such cases, the RL agent is more likely to encounter and select normal patches – especially in the early steps – leading to inefficient exploration and increased risk of a miss. The sparse distribution of tumor regions makes it difficult for the agent to consistently focus on diagnostically relevant areas within a limited sampling budget. A more efficient strategy would involve the early filtering of normal, redundant, or uninformative patches, allowing the agent to concentrate its computational efforts on a smaller, more informative subset of patches. This could improve both performance and inference efficiency.
Incorporation into clinical workflows - Our method is particularly suited for deployment in clinical settings, from two key standpoints – Explainability and Calibration. First, clinicians are likely to trust decisions made by AI agents only if these decisions can be readily and rationally explained. Our RL agent consistently selects high-attention, tumor-rich patches, as compared to normal patches. Moreover, this selection occurs far more frequently than that of a random policy. These selection strategies make the decisions of the RL agent intuitive and easy to explain to clinicians (Section˜4.5).
Second, clinicians are also likely to benefit from a reliable “confidence” score from the AI model, to know when they can, and cannot, trust its predictions. We showed that the expected calibration error (ECE) for SASHA=0.2 were consistently lower than those of competing sota attention-based models like ACMIL and DTFD (Section˜4.6). With well-calibrated models (lower ECE), the model’s confidence in its predictions (e.g., predictive uncertainty) is reliable. In this case, it becomes a useful indicator for when referral to a clinician or expert is warranted. In other words, by quantifying predictive uncertainty of the SASHA-0.2 classifier with established approaches (e.g., predictive entropy), our RL model could provide clinicians with valuable additional information about when its prediction can (or cannot) be trusted. Finally, inference time accuracy may depend on the extent of domain shift between the test and training clinical samples; developing approaches that combine domain adaptation with efficient sampling may be relevant to address this challenge.